OpenStudy (anonymous):

The two-way table shows the number of people from Group A or B who chose Option 1 or 2. What percent of people in Group A chose Option 2? Express your answer to the nearest whole percent. % Group A Group B Option 1 152 207 Option 2 106 95

OpenStudy (anonymous):

I can help you

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

@iGreen can you help real quik

OpenStudy (anonymous):

First, tell me how many people, in total, are in Group A? Then, how many people in Group A chose Option B?

OpenStudy (anonymous):

Option 2 **

OpenStudy (anonymous):

258=group A and 106 chose option 2

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

right, so now we get the percentage by: \(\sf \large \frac{ \text{Group A people who chose Option 2}}{\text{total number of people in GroupA}} \times 100%\)

OpenStudy (anonymous):

@Data_LG2 do i divide after multiplying?

OpenStudy (anonymous):

people who chose option 2 =10,600 total number =25,800

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

no no and no! :o Divide them first then multiply by 100. and you already told me the values that you should put in: \(\color{blue}{\text{Originally Posted by}}\) @NunChuckBoy3101 258=group A and 106 chose option 2 \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

T^T which part you did not understand? what I told you is pretty straightforward: right, so now we get the percentage by: \(\sf \large \frac{ \text{Group A people who chose Option 2}}{\text{total number of people in GroupA}} \times 100% =\frac{ 106}{258}\times 100%\)

OpenStudy (anonymous):

\(\sf \large \frac{ 106}{258} \times 100%\)

OpenStudy (anonymous):

ok so divide 106 and 258?

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@iGreen @ganeshie8

OpenStudy (anonymous):

243%?

OpenStudy (anonymous):

41% is correct

OpenStudy (anonymous):

@Data_LG2 is 41% correct

OpenStudy (anonymous):

106 divide it by 258 = 0.4108527131782 multiply the quotient by 100% = 41.08527131782 round the answer to the nearest whole number= 41%

OpenStudy (anonymous):

yes, that would be right :)

OpenStudy (anonymous):

ok thank you